### v1

@@ Abstract
Mapping the integrated 21cm emission line from dark matter-tracing neutral hydrogen gas is the primary science goal for MeerKLASS (MeerKAT’s Large Area Synoptic Survey). Prior to the arrival of MeerKAT, this intensity mapping technique had only been tested on a couple of pre-existing single-dish radio telescopes with a handful of observational hours with which to make early pioneering detections. The 64-dish MeerKAT array, precursor to the Square Kilometre Array Observatory (SKAO), can scan the sky in auto-correlation mode and perform intensity mapping across large sky areas, presenting the exciting potential for a wide-sky ( ${\gtrsim}\,10{,}000\,{\rm deg}^{2}$) spectroscopic survey across redshift 0.4<z< 1.450.4\,{<}\,z\,{<}\,1.45. Validating the auto-correlation (or single-dish) mode of observation for a multi-dish array and developing the analysis pipeline with which to make unbiased measurements has presented major challenges to this endeavour. In this work, we overview the advances in the field that have facilitated a robust analysis framework for single-dish intensity mapping, and review some results that showcase its success using early MeerKLASS surveys. We demonstrate our control of foreground cleaning, signal loss and map regridding to deliver detections of cosmological clustering within the intensity maps through cross-correlation power spectrum measurements with overlapping galaxy surveys. Finally, we discuss the prospects for future MeerKLASS observations and forecast its potential, making our code publicly available: https://github.com/meerklass/MeerFish.
@@ 1 Introduction
Mapping the large-scale structure of the Universe through the redshifted 21cm emission line of neutral hydrogen (Hi) is emerging as a powerful technique in observational cosmology. By surveying the diffuse Hi signal without resolving individual galaxies, Hi intensity mapping offers a means to efficiently chart matter fluctuations across vast cosmological areas and redshift ranges [31, 24, 154, 44, 93, 144].
@@
Prior to the arrival of MeerKAT111sarao.ac.za/science/meerkat/, 21cm intensity mapping had only been explored through a handful of pioneering efforts using existing single-dish radio telescopes such as the Green Bank Telescope (GBT) and Parkes, with restricted sky coverage and observational time [98, 137, 152, 13, 153]. These early detections demonstrated the feasibility of cross-correlation analyses between intensity maps and galaxy surveys, but enhanced sensitivity and a vast increase in observational resources would be needed to achieve competitive cosmological constraints. Surveys such as CHIME [12], FAST [92] and uGMRT [67], are aiming to exploit this opportunity, with more on the horizon [142, 51, 2]. The advent of MeerKAT, a 64-dish precursor to the Square Kilometre Array Observatory (SKAO)222skao.int/en, has also opened a new chapter in this field.
@@
MeerKAT’s Large Area Synoptic Survey (MeerKLASS)333meerklass.org/ [119] is a major effort to realise the scientific potential of Hi intensity mapping in the pre-SKAO era. While the primary operation of the MeerKAT array is as an interferometer, the distribution of its baselines is not well-suited to the very large angular scales required for Hi intensity mapping on “cosmological” scales ( ${\lesssim}\,0.2\,h\,\text{Mpc}^{-1}$) [36]. The solution is instead to use the auto-correlation of the voltage signal from each of the 64 dishes, effectively operating the array in single-dish mode to gather wide-field maps of the sky for each individual dish. This approach allows MeerKLASS to target the detection of the 21cm signal across thousands of square degrees, while at the same time retaining the interferometric visibilities for a wide range of commensal science. The frequency coverage from using MeerKAT’s UHF-band receiver will allow access to redshifts up to z= 1.45z\,{=}\,1.45 [SKAO Band 1 will extend this to z= 3z\,{=}\,3 128], making single-dish 21cm intensity mapping with these instruments ideally suited for vast volume spectroscopic surveys.
@@
Validating the single-dish mode of observation on a multi-dish array primarily designed for interferometry has presented significant technical and methodological challenges. Unlike traditional interferometric experiments, single-dish intensity mapping with MeerKAT has required a dedicated pipeline to calibrate, map-make, clean, and statistically analyse large volumes of novel data. Dominant foregrounds [117, 94, 149, 8], radio frequency interference (RFI) [75, 18, 68], correlated noise [91, 80], beam systematics [15, 99, 132, 48], and the inherently faint amplitude of the Hi brightness temperature, all compound the difficulty of extracting unbiased cosmological measurements.
@@
To meet these challenges, the MeerKLASS collaboration has developed a robust data treatment and analysis framework. Iterative self-calibration has been explored [145, 100] to suppress model reliance. Foreground cleaning methods, particularly blind component separation techniques [10, 40, 55, 132] such as Principal Component Analysis (PCA), are employed to reduce astrophysical foregrounds. Carefully accounting for signal loss from foreground cleaning is also crucial, and currently quantified and corrected using foreground transfer functions derived through mock signal injection into the real data [136, 57]. Another critical development has been the regridding of spherical sky maps into Cartesian volumes suitable for Fourier analysis, avoiding significant biases in large-sky surveys [53]. With these tools in hand, MeerKLASS has achieved statistically significant detections of Hi clustering through cross-correlations with overlapping spectroscopic galaxy surveys such as WiggleZ and GAMA [56, 100, 41]. These detections have marked a key milestone: the first robust demonstration of single-dish intensity mapping with a multi-dish array, laying the groundwork for full-scale cosmological analyses with the extended MeerKLASS and the upcoming SKAO.
@@ 2 Observations with MeerKLASS
Since 2018, MeerKLASS has been using the MeerKAT array [83] to steadily build survey area and sensitivity with single-dish mode observations. Initially, all observations were done using MeerKAT’s L-band, which has a receiver bandpass of $900\,{<}\,\nu\,{<}\,1670\,$MHz. However, since large cosmological volumes only become available for redshifted-21cm at  ${\lesssim}\,1200\,$MHz, plus RFI contamination further limits this window to $971\,{<}\,\nu\,{<}\,1023\,$MHz, the useable cosmological redshift coverage is limited to 0.39<z< 0.460.39\,{<}\,z\,{<}\,0.46. This partly motivated the move to the lower frequency receiver in UHF-band, which we discuss shortly. The first pilot MeerKLASS L-band survey was completed in 2019 [145], and then a deep-field survey in late 2021 [100]. This data served as essential proving grounds for developing calibration, map-making and analysis pipelines, and led to the first cosmological clustering detections with MeerKAT [56], much of which we will review.
@@
The current MeerKLASS approach is to conduct autocorrelation (single-dish) observations with a fast scanning strategy to suppress the impact of slow instrumental gain variations on the angular scales of interest for cosmology. The dishes scan back-and-forth in azimuth at constant elevation to minimise ground pickup and atmospheric fluctuations, with noise diodes firing at regular intervals to provide a stable reference for gain calibration. This approach ensures reasonably uniform sky coverage while maintaining instrumental stability. In parallel, the interferometric visibilities are also recorded in an on-the-fly (OTF) mode to support commensal science, and will be the subject of forthcoming publications [45, 97, 108]. We refer those readers interested in the technical implementation of this pipeline to Wang et al. [145], where the calibration, RFI flagging and map-making methodology was pioneered for this unique data (also see Li et al. [91], Irfan et al. [79], MeerKLASS Collaboration et al. [100] for related developments). The overview in this paper will not go into detail on this aspect and instead focuses on the post-calibration, analysis workflow, i.e. from maps to power spectra, as titled.
@@
MeerKLASS is now entering a mature phase. With large observational campaigns planned, it presents a fantastic opportunity for precision cosmology at radio wavelengths and complementarity with upcoming galaxy surveys such as DESI, Euclid, 4MOST and the Rubin Observatory [5, 33, 1, 96]. Table 1 summarises the cumulative MeerKLASS observations, both completed and planned. MeerKLASS is now steering towards using the UHF-band receiver in the most recent and future observations. As discussed, we found that the L-band receiver’s useful cosmological range was limited to 0.39<z< 0.460.39\,{<}\,z\,{<}\,0.46. The UHF-band however, offers a dramatically wider redshift window (0.4<z< 1.450.4\,{<}\,z\,{<}\,1.45), which early observations are showing is more robust to RFI contamination. The move to UHF holds crucial gains for cosmology since the observed comoving volume will be 40 times larger than an equivalent area of observation in the L-band.
@@
Year (Added) (Total) area Receiver Redshift Status
@@
2024 +270 380 1 600 UHF 0.4<z< 1.450.4\,{<}\,z\,{<}\,1.45 Observed
2025 +500 880 3 600 UHF 0.4<z< 1.450.4\,{<}\,z\,{<}\,1.45 Observing
2026 +500 1 380 5 600 UHF 0.4<z< 1.450.4\,{<}\,z\,{<}\,1.45 Awarded under XLP
2027 +550 1 930 7 800 UHF 0.4<z< 1.450.4\,{<}\,z\,{<}\,1.45 Extension under XLP
2028 +570 2 500 10 000 UHF 0.4<z< 1.450.4\,{<}\,z\,{<}\,1.45 Extension under XLP
2029 +500 3 000 12 000 UHF 0.4<z< 1.450.4\,{<}\,z\,{<}\,1.45 Pre-SKAO continuation
2030 +500 3 500 14 000 UHF 0.4<z< 1.450.4\,{<}\,z\,{<}\,1.45 Pre-SKAO continuation
@@ 3 Foreground cleaning
For an unbiased measurement of large-scale structure, these astrophysical foregrounds must be removed with high precision [14]. Because detailed foreground models remain uncertain, most successful approaches in current analyses adopt “blind” cleaning strategies that rely on the spectral smoothness of foregrounds rather than prior knowledge of their exact structure and amplitude [10, 40, 55, 132]. Principal Component Analysis (PCA)-based foreground cleaning has become one of the most widely used techniques [98, 13, 56, 41, 100, 48], exploiting statistical decompositions of the data to identify and subtract dominant smooth modes while preserving the cosmological signal.
@@ 3.1.1 Multiscale extension (mPCA)
Instead of relying only on the radial characterisation of the contaminants, it is also helpful to consider the pixel-domain (angular) information. The rationale is that the systematic influence over astrophysical foreground happens differently at different spatial angular scales, being more prominent at the small scales. Instead, at large scales, the astrophysical contribution remains the most dominant. We have put this idea into practice with MeerKLASS data [41], designing a multiscale PCA (mPCA) algorithm that we compared to standard PCA.
In this multiscale approach, each frequency map is first decomposed into wavelet-filtered large- and small-scale maps, using the isotropic undecimated wavelet transform [134]. From one data cube, we now have two cubes of the same size (each of them referring to the large- and small-scale only information) which are then cleaned separately: the contaminant separation problem becomes two independent problems, with their own, independently defined mixing matrix and number of removed components ($N_{\mathrm{fg}}$). This allows for a better characterisation of the two contaminant ‘regimes’. The final cleaned map is the sum of the independently cleaned small- and large-scale maps, since the wavelet decomposition is exact.
By using one of our cross-correlation detections as a benchmark [56], we could test the performance of mPCA. We found mPCA led to a higher and more scale-independent cross-correlation signal, plus evidence for less residual contamination. The mPCA maps also had less prominent off-diagonal terms in its frequency-correlation matrix, suggesting less mode-mixing between residual contaminants and signal, making it more robust. We refer to Carucci et al. [41] for more details. In summary, mPCA offers a promising extension for future Hi intensity mapping surveys.
@@ 3.2 Signal loss correction: the foreground transfer function
Blind foreground cleaning has shown impressive progress in removing the overwhelming astrophysical contaminants, enabling the recovery of the cosmological Hi brightness fluctuations. However, since the cleaning process projects out correlated modes along the line of sight, a fraction of the true Hi signal will inevitably be removed as well [136, 49]. This effect, known as signal loss, is unavoidable and must be carefully corrected for.
Whilst considerable effort has been devoted to mitigating the bias from foreground residuals, the equally important problem of over-cleaning is less often addressed. Over-cleaning leads to signal loss that can substantially suppress the measured power spectrum, particularly on large radial scales most degenerate with foregrounds. Unlike residual contamination, the bias from signal loss is not mitigated by cross-correlations, thus arguably presenting a more threatening problem. The MeerKLASS collaboration has therefore carried out dedicated studies of signal loss, with the aim of ensuring robust cosmological interpretation of foreground-cleaned maps [57].
@@
In practice, $\widehat{\mathcal{T}}(\bm{k})$ is calculated pseudo-empirically by injecting mock Hi signals into the real contaminated data, applying the same cleaning procedure, and measuring the fractional suppression of the mock power spectrum [136]. The reason for using the actual data itself is that accurate emulation of the foregrounds and their response to the complex instrumentation is extremely challenging and would inevitably result in underestimation of the contamination.
Construction recipe: Following Cunnington et al. [57], the recipe for building an unbiased transfer function proceeds as:
@@
This empirical approach accounts for the complex correlations and anti-correlations between signal and foreground modes, which would otherwise be neglected, leading to under-estimation of signal loss [as demonstrated in 57]. The subtraction of the square-brackets term (i.e. the foreground-cleaned data) in Equation 7 can be useful for reducing variance in the transfer function. However, further utility from the transfer function is to use it as a basis for covariance estimation (discussed in Section 5). In this case, it is important that the square bracket term not be subtracted so as not to under-estimate error bars.
Work from the MeerKLASS collaboration has shown that even aggressive cleaning (removing many modes) can be reliably corrected. The left panel of Figure 1 shows the reconstructed spherically averaged power spectrum is unbiased at the sub-percent level on intermediate and small scales, and within uncertainties on the largest scales. Here, a set of $N$-body semi-analytical simulations [see 57, for full details] was treated as the real “observations” with true power $P_{\textrm{H{i}}}$. Simulated foreground contamination was added to this and cleaned using standard PCA. Then, simpler lognormal mocks were used to construct the transfer function $\mathcal{T}$. The known truth of the simulated data allows us to validate the ability of the transfer function to reconstruct the signal by comparing it to the original, foreground-free, power $P_{\textrm{H{i}}}$. This was tested for two cases with $N_{\text{fg}}\,{=}\,8$ and 12 PCA modes removed. Similar accuracy is achieved in both cases, with a slightly larger variance for $N_{\text{fg}}\,{=}\,12$ due to the more aggressive cleaning.
@@
Application in precision cosmology: A surprising and beneficial property of the transfer function is its weak dependence on the assumed cosmological model used in the injected mock simulations. In principle, the construction of the transfer function requires a fiducial cosmology to generate the input Hi power spectrum, and a mismatch between the assumed and true cosmology could in turn bias the recovered clustering statistics. However, it was empirically demonstrated in Cunnington et al. [57] that such discrepancies yield only minimal errors in parameter inference. This robustness is particularly encouraging for ultra-large scale probes such as local primordial non-Gaussianity, $f_{\mathrm{NL}}$ [23], where the largest modes most affected by signal loss are also the most important for improved constraints.
The origin of this robustness was clarified in Chen [47], which showed using a quadratic estimator formalism that the transfer function is mathematically equivalent to a normalisation of the window function of the estimator, and therefore depends only on the linear operator that performs the foreground cleaning. In this picture, the transfer function can be interpreted as a renormalisation that also encapsulates mode-mixing induced by PCA cleaning, and is thus independent of the cosmology of the mock signal. The work in Chen [47] also provided a formal proof that, owing to the orthogonality of the PCA eigenmodes, the correction for both cross- and auto-power is $\mathcal{T}^{-1}(\bm{k})$, contrary to the previously assumed $\mathcal{T}^{-2}$ for the auto-spectrum.
Chen [47] also highlighted the important role mode-mixing (caused by the PCA) can have on the signal reconstruction. This revealed that the transfer function in its form defined by Equation 8 and Equation 7, will include effects of mode-mixing, not just signal loss. To isolate the pure signal loss, and enforce further model independence, the alternative approach, where the eigenmodes are not recalculated with mock injection, can be taken, i.e. using the below in Equation 8 instead:
@@
Without an additional correction for the mode mixing, this approach actually delivers a biased power spectra recovery [shown in the upper panel of Figure C1 of 57]. However, as motivated by Chen [47], it provides a more solid foundation to treat the complex effects of mode-mixing separately from signal loss. Hence, future MeerKLASS work will pursue this route for robust, model-independent signal loss correction, with increased suitability for precision cosmology.

### v2

@@ Abstract
Mapping the integrated 21cm emission line from dark matter-tracing neutral hydrogen gas is the primary science goal for MeerKLASS (MeerKAT’s Large Area Synoptic Survey). Prior to the arrival of MeerKAT, this intensity mapping technique had only been tested on a couple of pre-existing single-dish radio telescopes with a handful of observational hours with which to make early pioneering detections. The 64-dish MeerKAT array, precursor to the SKA Observatory (SKAO), can scan the sky in auto-correlation (or single-dish) mode and perform intensity mapping across large sky areas, presenting the exciting potential for a wide area ( ${\gtrsim}\,10{,}000\,{\rm deg}^{2}$) spectroscopic survey across redshift 0.4<z< 1.450.4\,{<}\,z\,{<}\,1.45. Validating the single-dish mode of observation for a multi-dish array and developing the analysis pipeline with which to make unbiased measurements has presented major challenges to this endeavour. In this work, we overview the advances in the field that have facilitated a robust analysis framework for single-dish intensity mapping, and review some results that showcase its success using early MeerKLASS surveys. We demonstrate our control of foreground cleaning, signal loss and map regridding to deliver detections of cosmological clustering within the intensity maps through cross-correlation power spectrum measurements with overlapping galaxy surveys. Finally, we discuss the prospects for future MeerKLASS observations and forecast its potential, making our code publicly available: https://github.com/meerklass/MeerFish.
@@ 1 Introduction
Mapping the large-scale structure of the Universe through the redshifted 21cm emission line of neutral hydrogen (Hi) is emerging as a powerful technique in observational cosmology. By surveying the diffuse Hi signal without resolving individual galaxies, Hi intensity mapping offers a means to efficiently chart matter fluctuations across vast cosmological areas and redshift ranges [Bharadwaj:2000av, Battye:2004re, Wyithe:2007rq, Chang:2007xk, Liu:2019awk, Villaescusa-Navarro:2018vsg].
@@
Prior to the arrival of MeerKAT111sarao.ac.za/science/meerkat/, CHIME [CHIME:2022dwe], FAST [Hu:2019okh] and uGMRT [Elahi:2024qbp], 21cm intensity mapping had only been explored through a handful of pioneering efforts using existing single-dish radio telescopes such as the Green Bank Telescope (GBT) and Parkes, with restricted sky coverage and observational time [Masui:2012zc, Switzer:2013ewa, Wolz:2015lwa, Anderson:2017ert, eBOSS:2021ebm]. These early detections demonstrated the feasibility of cross-correlation analyses between intensity maps and galaxy surveys, but enhanced sensitivity and a vast increase in observational resources would be needed to achieve competitive cosmological constraints.
[…]
@@
MeerKAT’s Large Area Synoptic Survey (MeerKLASS)333meerklass.org/ [MeerKLASS:2017vgf] is a major effort to realise the scientific potential of Hi intensity mapping in the pre-SKAO era. While the primary operation of the MeerKAT array is as an interferometer, the distribution of its baselines is not well-suited to the very large angular scales required for Hi intensity mapping on “cosmological” scales ( ${\lesssim}\,0.2\,h\,\text{Mpc}^{-1}$) [Bull:2014rha]. The solution is instead to use the auto-correlation of the voltage signal from each of the 64 dishes, effectively operating the array in single-dish mode [^4] From hereon we will exclusively refer to the dish auto-correlation mode of observations as “single-dish mode” to help distinguish it from the different concept of auto-correlation (Hi-Hi) power spectrum measurements, discussed later. to gather wide-field maps of the sky for each individual dish. This approach allows MeerKLASS to target the detection of the 21cm signal across thousands of square degrees, while at the same time retaining the interferometric visibilities for a wide range of commensal science. The frequency coverage from using MeerKAT’s UHF-band receiver will allow access to redshifts up to z= 1.45z\,{=}\,1.45 [SKAO Band 1 will extend this to z= 3z\,{=}\,3 SKA:2018ckk], making single-dish 21cm intensity mapping with these instruments ideally suited for vast volume spectroscopic surveys.
@@
Validating the single-dish mode of observation on a multi-dish array primarily designed for interferometry has presented significant technical and methodological challenges. Unlike traditional interferometric experiments, single-dish intensity mapping with MeerKAT has required a dedicated pipeline to calibrate, map-make, clean, and statistically analyse large volumes of novel data. Dominant foregrounds [Santos:2004ju, Liu:2009qga, Wolz:2013wna, Alonso:2014sna], radio frequency interference (RFI) [Harper:2018ncl, 2019JAI.....840010B, Engelbrecht:2024eoc], correlated noise [Li:2020bcr, Irfan:2023njr], beam systematics [2021MNRAS.502.2970A, Matshawule:2020fjz, Spinelli:2021emp, Chen:2025bcx], and the inherently faint amplitude of the Hi brightness temperature, all compound the difficulty of extracting unbiased cosmological measurements.
@@
To meet these challenges, the MeerKLASS collaboration has developed a robust data treatment and analysis framework. Iterative self-calibration has been explored [Wang:2020lkn, MeerKLASS:2024ypg] to suppress model reliance. Foreground cleaning methods, particularly blind component separation techniques [Alonso:2014dhk, Carucci:2020enz, Cunnington:2020njn, Spinelli:2021emp] such as Principal Component Analysis (PCA), are employed to reduce astrophysical foregrounds. Carefully accounting for signal loss from foreground cleaning is also crucial, and currently quantified and corrected using foreground transfer functions derived through mock signal injection into the real data [Switzer:2015ria, Cunnington:2023jpq]. Another critical development has been the regridding of spherical sky maps into Cartesian volumes suitable for Fourier analysis, avoiding significant biases in large-sky surveys [Cunnington:2023aou]. With these tools in hand, MeerKLASS has achieved statistically significant detections of Hi clustering through cross-correlations with overlapping spectroscopic galaxy surveys such as WiggleZ and GAMA [Cunnington:2022uzo, MeerKLASS:2024ypg, Carucci:2024qpm]. These detections have marked a key milestone: the first robust demonstration of single-dish intensity mapping with a multi-dish array, laying the groundwork for full-scale cosmological analyses with the extended MeerKLASS and the upcoming SKAO.
@@ 2 Observations with MeerKLASS
Since 2018, MeerKLASS has been using the MeerKAT array [2016mks..confE...1J] to steadily build survey area and sensitivity with single-dish mode observations. Initially, all observations were done using MeerKAT’s L-band, which has a receiver bandpass of $900\,{<}\,\nu\,{<}\,1670\,$MHz. However, since large cosmological volumes only become available for redshifted-21cm at  ${\lesssim}\,1200\,$MHz, plus RFI contamination further limits this window to $971\,{<}\,\nu\,{<}\,1023\,$MHz, the useable cosmological redshift coverage is limited to 0.39<z< 0.460.39\,{<}\,z\,{<}\,0.46. This partly motivated the move to the lower frequency receiver in UHF-band, which we discuss shortly. The first pilot MeerKLASS L-band survey was completed in 2019 [Wang:2020lkn], and then a deep-field survey in late 2021 [MeerKLASS:2024ypg]. This data served as essential proving grounds for developing calibration, map-making and analysis pipelines, and led to the first cosmological clustering detections with MeerKAT [Cunnington:2022uzo], much of which we will review.
@@
The current MeerKLASS approach is to conduct single-dish observations with a fast scanning strategy to suppress the impact of slow instrumental gain variations on the angular scales of interest for cosmology. The dishes scan back-and-forth in azimuth at constant elevation to minimise ground pickup and atmospheric fluctuations, with noise diodes firing at regular intervals to provide a stable reference for gain calibration. This approach ensures reasonably uniform sky coverage while maintaining instrumental stability. In parallel, the interferometric visibilities are also recorded in an on-the-fly (OTF) mode to support commensal science, and will be the subject of forthcoming publications [ChatterjeeOTF, ManglaOTF, PaulOTF]. We refer those readers interested in the technical implementation of this pipeline to Wang:2020lkn, where the calibration, RFI flagging and map-making methodology was pioneered for this unique data (also see Li:2020bcr, Irfan:2021xuk, MeerKLASS:2024ypg for related developments). The overview in this paper will not go into detail on this aspect and instead focuses on the post-calibration, analysis workflow, i.e. from maps to power spectra, as titled.
@@
MeerKLASS is now entering a mature phase. With large observational campaigns planned, it presents a fantastic opportunity for precision cosmology at radio wavelengths and complementarity with upcoming galaxy surveys such as DESI, Euclid, 4MOST and the Rubin Observatory [DESI:2016fyo, Euclid:2019clj, 2019Msngr.175....3D, LSSTDarkEnergyScience:2018jkl]. Table 1 summarises the cumulative MeerKLASS observations, both completed and planned. MeerKLASS is now steering towards using the UHF-band receiver in the most recent and future observations. As discussed, we found that the L-band receiver’s useful cosmological range was limited to 0.39<z< 0.460.39\,{<}\,z\,{<}\,0.46. The UHF-band however, offers a dramatically wider redshift window (0.4<z< 1.450.4\,{<}\,z\,{<}\,1.45), which early observations are showing is more robust to RFI contamination. The move to UHF holds crucial gains for cosmology since the observed comoving volume will be 40 times larger than an equivalent area of observation in the L-band.
@@
Year (Added) (Total) area Receiver Redshift Status as of 2026
@@
2024 +270 380 1,600 UHF 0.4<z< 1.450.4\,{<}\,z\,{<}\,1.45 Observed
2025 +500 880 3,600 UHF 0.4<z< 1.450.4\,{<}\,z\,{<}\,1.45 Observing
2026 +500 1,380 5,600 UHF 0.4<z< 1.450.4\,{<}\,z\,{<}\,1.45 Awarded under XLP
2027 +550 1,930 7,800 UHF 0.4<z< 1.450.4\,{<}\,z\,{<}\,1.45 Extension under XLP
2028 +570 2,500 10,000 UHF 0.4<z< 1.450.4\,{<}\,z\,{<}\,1.45 Extension under XLP
2029 +500 3,000 12,000 UHF 0.4<z< 1.450.4\,{<}\,z\,{<}\,1.45 Pre-SKAO continuation
2030 +500 3,500 14,000 UHF 0.4<z< 1.450.4\,{<}\,z\,{<}\,1.45 Pre-SKAO continuation
@@ 3 Foreground cleaning
For an unbiased measurement of large-scale structure, these astrophysical foregrounds must be removed with high precision [Ansari:2011bv]. Because detailed foreground models remain uncertain, most successful approaches in current analyses adopt “blind” cleaning strategies that rely on the spectral smoothness of foregrounds rather than prior knowledge of their exact structure and amplitude [Alonso:2014dhk, Carucci:2020enz, Cunnington:2020njn, Spinelli:2021emp]. Principal Component Analysis (PCA)-based foreground cleaning has become one of the most widely used techniques [Masui:2012zc, Anderson:2017ert, Cunnington:2022uzo, Carucci:2024qpm, MeerKLASS:2024ypg, Chen:2025bcx], exploiting statistical decompositions of the data to identify and subtract dominant smooth modes while preserving the cosmological signal.
@@ 3.1.1 Multiscale extension (mPCA)
Instead of relying only on the radial characterisation of the contaminants, it is also helpful to consider the pixel-domain (angular) information. The rationale is that the systematic influence over astrophysical foreground happens differently at different spatial angular scales, being more prominent at the small scales. Instead, at large scales, the astrophysical contribution remains the most dominant. We have put this idea into practice with MeerKLASS data [Carucci:2024qpm], designing a multiscale PCA (mPCA) algorithm that we compared to standard PCA.
In this multiscale approach, each frequency map is first decomposed into wavelet-filtered large- and small-scale maps, using the isotropic undecimated wavelet transform [Starck2007]. From one data cube, we now have two cubes of the same size (each of them referring to the large- and small-scale only information) which are then cleaned separately: the contaminant separation problem becomes two independent problems, with their own, independently defined mixing matrix and number of removed components ($N_{\mathrm{fg}}$). This allows for a better characterisation of the two contaminant ‘regimes’. The final cleaned map is the sum of the independently cleaned small- and large-scale maps, since the wavelet decomposition is exact.
By using one of our cross-correlation detections as a benchmark [Cunnington:2022uzo], we could test the performance of mPCA. We found mPCA led to a higher and more scale-independent cross-correlation signal, plus evidence for less residual contamination. The mPCA maps also had less prominent off-diagonal terms in its frequency-correlation matrix, suggesting less mode-mixing between residual contaminants and signal, making it more robust. We refer to Carucci:2024qpm for more details. In summary, mPCA offers a promising extension for future Hi intensity mapping surveys.
@@ 3.2 Signal loss correction: the foreground transfer function
Blind foreground cleaning has shown impressive progress in removing the overwhelming astrophysical contaminants, enabling the recovery of the cosmological Hi brightness fluctuations. However, since the cleaning process projects out correlated modes along the line of sight, a fraction of the true Hi signal will inevitably be removed as well [Switzer:2015ria, Cheng:2018osq]. This effect, known as signal loss, is unavoidable and must be carefully corrected for.
Whilst considerable effort has been devoted to mitigating the bias from foreground residuals, the equally important problem of over-cleaning is less often addressed. Over-cleaning leads to signal loss that can substantially suppress the measured power spectrum, particularly on large radial scales most degenerate with foregrounds. Unlike residual contamination, the bias from signal loss is not mitigated by cross-correlations, thus arguably presenting a more threatening problem. The MeerKLASS collaboration has therefore carried out dedicated studies of signal loss, with the aim of ensuring robust cosmological interpretation of foreground-cleaned maps [Cunnington:2023jpq].
@@
In practice, $\widehat{\mathcal{T}}(\bm{k})$ is calculated pseudo-empirically by injecting mock Hi signals into the real contaminated data, applying the same cleaning procedure, and measuring the fractional suppression of the mock power spectrum [Switzer:2015ria]. The reason for using the actual data itself is that accurate emulation of the foregrounds and their response to the complex instrumentation is extremely challenging and would inevitably result in underestimation of the contamination.
The transfer function correction is only empirically accessible through mock signal injection, meaning its accuracy ultimately relies on the assumption that the injected mocks adequately sample the relevant signal–foreground couplings present in the real data. Large numbers of mock simulations  ${\gtrsim}\,100$ are typically required to reach a converged transfer function, making this a computationally expensive process. Furthermore, as recently clarified by [Chen:2025til], blind cleaning induces mode-mixing in Fourier space, which will propagate into the transfer function calculation. We discuss some of the consequences of this at the end of this section.
Construction recipe: Following Cunnington:2023jpq, the recipe for building an unbiased transfer function proceeds as:
@@
This empirical approach accounts for the complex correlations and anti-correlations between signal and foreground modes, which would otherwise be neglected, leading to under-estimation of signal loss [as demonstrated in Cunnington:2023jpq]. The subtraction of the square-brackets term (i.e. the foreground-cleaned data) in Equation 7 can be useful for reducing variance in the transfer function. However, further utility from the transfer function is to use it as a basis for covariance estimation (discussed in Section 5). In this case, it is important that the square bracket term not be subtracted so as not to under-estimate error bars.
Work from the MeerKLASS collaboration has shown that even aggressive cleaning (removing many modes) can be reliably corrected. The left panel of Figure 1 shows the reconstructed spherically averaged power spectrum is unbiased at the sub-percent level on intermediate and small scales, and within uncertainties on the largest scales. Here, a set of $N$-body semi-analytical simulations [see Cunnington:2023jpq, for full details] was treated as the real “observations” with true power $P_{\textrm{H{i}}}$. Simulated foreground contamination was added, including diffuse Galactic synchrotron and free-free emission, bright point sources, along with some instrumental polarisation leakage to increase the spectral complexity. This foreground contaminated simulation was then cleaned using standard PCA, and then simpler lognormal Hi-mocks were used to construct the transfer function $\mathcal{T}$. The known truth of the simulated data allows us to validate the ability of the transfer function to reconstruct the signal by comparing it to the original, foreground-free, power $P_{\textrm{H{i}}}$. This was tested for two cases with $N_{\text{fg}}\,{=}\,8$ and 12 PCA modes removed. Similar accuracy is achieved in both cases, with a slightly larger variance for $N_{\text{fg}}\,{=}\,12$ due to the more aggressive cleaning.
@@
Application in precision cosmology: A surprising and beneficial property of the transfer function is its weak dependence on the assumed cosmological model used in the injected mock simulations. In principle, the construction of the transfer function requires a fiducial cosmology to generate the input Hi power spectrum, and a mismatch between the assumed and true cosmology could in turn bias the recovered clustering statistics. However, it was empirically demonstrated in Cunnington:2023jpq that such discrepancies yield only minimal errors in parameter inference. This robustness is particularly encouraging for ultra-large scale probes such as local primordial non-Gaussianity, $f_{\mathrm{NL}}$ [Bartolo:2004if], where the largest modes most affected by signal loss are also the most important for improved constraints.
The origin of this robustness was clarified in Chen:2025til, which showed using a quadratic estimator formalism that the transfer function is mathematically equivalent to a normalisation of the window function of the estimator, and therefore depends only on the linear operator that performs the foreground cleaning. In this picture, the transfer function can be interpreted as a renormalisation that also encapsulates mode-mixing induced by PCA cleaning, and is thus independent of the cosmology of the mock signal. The work in Chen:2025til also provided a formal proof that, owing to the orthogonality of the PCA eigenmodes, the correction for both cross- and auto-power is $\mathcal{T}^{-1}(\bm{k})$, contrary to the previously assumed $\mathcal{T}^{-2}$ for the auto-spectrum.
Chen:2025til also highlighted the important role mode-mixing (caused by the PCA) can have on the signal reconstruction. This revealed that the transfer function in its form defined by Equation 8 and Equation 7, will include effects of mode-mixing, not just signal loss. To isolate the pure signal loss, and enforce further model independence, an alternative approach, where the eigenmodes are not recalculated with mock injection, can be taken, i.e. using the below in Equation 8 instead:
@@
Without an additional correction for the mode mixing, this approach actually delivers a biased power spectra recovery [shown in the upper panel of Figure C1 of Cunnington:2023jpq]. However, as motivated by Chen:2025til, it provides a more solid foundation to treat the complex effects of mode-mixing separately from signal loss. Hence, future MeerKLASS work will pursue this route for robust, model-independent signal loss correction, with increased suitability for precision cosmology.